OpenStudy (lizilizi28):

simplify the expression if is possible log . 0.00001

OpenStudy (anonymous):

\[.00001=1\times10^{-5}\\ \log a^m=m\log a\]

OpenStudy (anonymous):

put 2-2 together :)

OpenStudy (lizilizi28):

what u mean 2-2

OpenStudy (lizilizi28):

i think the aswer is -5 im correct ?

Directrix (directrix):

I got -5.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

assuming "log" means base 10, o'course